OpenStudy (anonymous):

Factor completely 16y^2+25

OpenStudy (ash2326):

@CAT2142276 have you studied complex numbers? or Is this your question? \[16y^2-25\]

OpenStudy (anonymous):

yes that is my question , i dont know how to write the little 2 with out the carrot button..

OpenStudy (ash2326):

To write like that use equation:)

OpenStudy (ash2326):

Ok let's factor \[16y^2-25\] We know \[16=4^2\] and \(25=5^2\) so \[4^2y^2-5^2\] This can be written as \[(4y)^2-5^2\] Now we know \[(a^2-b^2)=(a-b)(a+b)\] Using this here we get \[\huge (4y)^2-(5)^2=(4y+5)(4y-5)\]

OpenStudy (anonymous):

oh okay thank you ;-)

OpenStudy (ash2326):

@CAT2142276 Welcome to Open Study:D

OpenStudy (anonymous):

how about this \[4a^{2}-16a+16\]

OpenStudy (ash2326):

@CAT2142276 I can help you here but if you close this question and post this one as new question. You'd get more users to help. Willy you do this?

OpenStudy (anonymous):

okay